# Entanglement and particle production from cosmological perturbations: a quantum optical simulation approach

### v1

@@ 1 Introduction
with $|\alpha_{k}|^{2}-|\beta_{k}|^{2}=1$. A nonzero $\beta_{k}$ means that what started out as a vacuum state at early times ends up containing particles when viewed at late times, with each mode having an average particle number of $|\beta_{k}|^{2}$ mukhanov1992theory ; parker1969quantized .
@@
The hallmark of the two-mode squeezed state is that the perturbed modes $k$ and k-k are highly entangled. The strong correlations between the entangled modes imply that the reduced density matrix of the squeezed state, which is mixed through a subtle coarse-graining, gives the non-vanishing von Neumann entropy. In brahma2020entanglement the coarse-graining entropy was done by averaging over the squeezing angle, which provides a stochastic part that lead the reduced density matrix’s off diagonal elements to zero. This process gives rise to entanglement entropy, which can account for part of the large entropy observed in the post-inflationary universe gasperini1993entropy ; boutivas2023entanglement . Also, many of the papers have argued that the squeezing phenomenon in the context of cosmology is associated with the emergence of classical behavior of our early quantum universe grishchuk1990squeezed ; albrecht1994inflation ; polarski1996semiclassicality ; lesgourgues1997quantum ; kiefer1998quantum ; kiefer2009quantum ; perez2006quantum ; brandenberger1990classical ; calzetta1995quantum ; Li:2024ljz ; Li:2024iji .
@@
This work contributes to the growing efforts of applying quantum optics and quantum information theory to the study of inflationary cosmology. Recent studies have demonstrated that quantum optical methods like continuous-variable quantum circuits, Gaussian-state manipulation, and entanglement measures provide powerful frameworks to explore primordial quantum states Liu_2021 ; Boutivas_2024 . Several works have explored simulating inflationary cosmology using quantum tools. Digital simulation of particle creation in expanding universes were done using quantum circuits and extended field theory algorithms Lloyd:2013xba ; Bao:2017iye ; Yang_2020 ; Maceda_2025 ; Piotrak:2025zhy . Earlier studies showed that quantum fluctuations during inflation naturally form squeezed states parker1969quantized ; grishchuk1990quantum . Analog models like Bose-Einstein condensates and coupled mechanical oscillators have also been used to mimic inflation-like behaviour Fischer_2004 ; Rhyno_2025 . Continuous variable photonic systems have been proposed to simulate quantum fields using the Gaussian formalism Marshall_2015 . Other works studied entropy and decoherence, linking squeezing to the emergence of classical structures grishchuk1990quantum ; albrecht1994inflationary ; Zhai:2024odw . Together, these works provide valuable context and motivation for our approach.

### v2

@@ 1 Introduction
with $|\alpha_{k}|^{2}-|\beta_{k}|^{2}=1$. A nonzero $\beta_{k}$ means that what started out as a vacuum state at early times ends up containing particles when viewed at late times, with each mode having an average particle number of $|\beta_{k}|^{2}$ mukhanov1992theory ; parker1969quantized ; jiang2025cosmological .
@@
The hallmark of the two-mode squeezed state is that the perturbed modes $k$ and k-k are highly entangled. The strong correlations between the entangled modes imply that the reduced density matrix of the squeezed state, which is mixed through a subtle coarse-graining, gives the non-vanishing von Neumann entropy. In brahma2020entanglement the coarse-graining entropy was done by averaging over the squeezing angle, which provides a stochastic part that lead the reduced density matrix’s off diagonal elements to zero. This process gives rise to entanglement entropy, which can account for part of the large entropy observed in the post-inflationary universe gasperini1993entropy ; boutivas2023entanglement ; martin2021real ; Martin:2021xml . Also, many of the papers have argued that the squeezing phenomenon in the context of cosmology is associated with the emergence of classical behavior of our early quantum universe grishchuk1990squeezed ; albrecht1994inflation ; polarski1996semiclassicality ; lesgourgues1997quantum ; kiefer1998quantum ; kiefer2009quantum ; perez2006quantum ; brandenberger1990classical ; calzetta1995quantum ; Li:2024ljz ; Li:2024iji ; colas2022four .
@@
This work contributes to the growing efforts of applying quantum optics and quantum information theory to the study of inflationary cosmology. Recent studies have demonstrated that quantum optical methods like continuous-variable quantum circuits, Gaussian-state manipulation, and entanglement measures provide powerful frameworks to explore primordial quantum states Liu_2021 ; Boutivas_2024 . Several works have explored simulating inflationary cosmology using quantum tools. For example, digital simulation of particle creation in expanding universes were done using quantum circuits and extended field theory algorithms Lloyd:2013xba ; Bao:2017iye ; Yang_2020 ; Maceda_2025 ; Piotrak:2025zhy . The redistribution of Gaussian entanglement for continuous variable of scalar field in an expanding universe was also analyzed by modeling the expansion as a Gaussian channel Li2022GaussianEntanglement . Earlier studies showed that quantum fluctuations during inflation naturally form squeezed states parker1969quantized ; grishchuk1990quantum . Continuous variable photonic systems have been proposed to simulate quantum fields using the Gaussian formalism Marshall_2015 . Other works studied entropy and decoherence, linking squeezing to the emergence of classical structures grishchuk1990quantum ; albrecht1994inflationary ; Zhai:2024odw . Together, these works provide valuable context and motivation for our approach.
In addition to quantum-optical simulations, condensed-matter systems have long been studied as analog models of cosmology. Analog models like Bose-Einstein condensates and coupled mechanical oscillators have been used as laboratory analogues of the early universe mimicking inflation-like expansion and particle pair production. Fischer_2004 ; Rhyno_2025 ; PhysRevA.103.023322 ; PhysRevD.110.125023 ; PhysRevLett.123.180502 . Experiment with the quantum fluids of light have been shown that a sudden change in a 33D photon fluid can cause spontaneous creation of particles, forming acoustic-peak patterns like those cosmic inflation Steinhauer_2022 . Various systems from expanding superfluid rings to photonic condensates and optical fiber circuits have been explored as analogues for early universe studies Pal_2024 . Other works include probing sonic analogues of Gibbons–Hawking radiation, trans-Planckian effects, and quantum backreaction, and exploring phenomena such as analogue gravitational lensing, black hole information scrambling, nonlinear quantum effects, and phase transitions in dipolar systems Fedichev2003PRL ; Fischer2004PRA ; Fischer2022EPJC ; Fischer2022PRD ; Fedichev2004PRD ; Fischer2017PRL ; Fischer2022PRA ; Fischer2023CQG ; Fischer2023PRD ; Fischer2024PRD ; Fischer2025arxiv ; Visser2002PRL . Besides quantum-optical and condensed-matter settings, synthetic mechanical lattices have also been employed to study analogues of inflationary and FLRW cosmology. In Refs. rhyno2025mechanical it is demonstrated how the bosonic dynamics of scalar-field fluctuations are mapped onto the motion of mechanical oscillators, and key inflationary quantities such as particle production and the power spectrum are simulated and measured. These above approaches from the divergent field in contrast to circuit based method we employed, creates the motivation to explore more on this subject matter and discover the unknown.